OpenStudy (imnotgoodatmath):

Which statement describes the error made in the following solution process? k minus seven equals negative eight The addition calculation was incorrect. Subtraction should have been used instead of addition. Seven was added to the left side of the equation instead of eight. Eight was added to the right side of the equation instead of seven.

OpenStudy (turingtest):

you appear to be missing some info

OpenStudy (mathstudent55):

Where is the solution we need to check?

OpenStudy (imnotgoodatmath):

oh, sorry ill type it again

OpenStudy (mathstudent55):

Can you please use numbers instead of spelling out numbers?

OpenStudy (imnotgoodatmath):

k - 7 = -8 +7 +8 k = 0

OpenStudy (igreen):

D..

OpenStudy (igreen):

You're supposed to add 7 to both sides of the equation..not 8.

OpenStudy (mathstudent55):

When you solve an equation, you are allowed to do the same to both sides of the equation. For example, you can add the same number to both sides. You can't add different numbers to both sides.

OpenStudy (imnotgoodatmath):

ok

OpenStudy (igreen):

k - 7 = -8 k - 7 + 7 = -8 + 7 k = -1

OpenStudy (imnotgoodatmath):

thx again iGreen :D

OpenStudy (igreen):

No problem :)